OpenStudy (anonymous):

An elevator car whose floor to ceiling distance is equal to 2.7m , starts from rest and ascends with a constant acceleration of 1.2m/s^2 .Two seconds Later a bolt drops from the cars ceiling the time taken by the bolt to hit the floor is?

OpenStudy (anonymous):

The answer should be 0.7s

OpenStudy (anonymous):

The actual speed at the starting point does not matter. This is because the bolts relative speed to the car is zero. Now the cars acceleration + the gravitational acceleration will give the bolt's acceleration rate relative to the cars. Start off like this: \[a = g + 1.2m/ s^{2} = 1.2m/s^{2} \] \[s = a *t ^{2} /2\] \[t =\sqrt{2s/a} = 0.69\approx0.7\] Hope that helps! n_n

OpenStudy (anonymous):

Thaxxx a lot@Itz_U_Girl

OpenStudy (anonymous):

Your welcome @open_study1 n_n